OpenStudy (anonymous):

A line has a slope of −23 and goes through the point (−5,2). If the point (37,b) lies on the line, what is the value of b?

OpenStudy (anonymous):

Remember the equation for the slope \(m\) given two points \((x_1,y_1),(x_2,y_2)\)\[ m=\frac{y_2-y_1}{x_2-x_1} \]

OpenStudy (anonymous):

In this case\[ -23 = \frac{b-2}{37-(-5)} \]

OpenStudy (anonymous):

You need to solve for \(b\)

OpenStudy (anonymous):

so the bootom equals 42?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so whats the answer buddy (:

OpenStudy (anonymous):

You need to solve for \(b\).

OpenStudy (anonymous):

\[ -23=\frac{b-2}{42} \]

OpenStudy (anonymous):

What do you do next to isolate the variable?

OpenStudy (anonymous):

so its -966?

OpenStudy (anonymous):

-968

OpenStudy (zale101):

i got -966

OpenStudy (zale101):

we had to do cross multiply